OpenStudy (kbreaux9):

For homework, Fred has to read 30 minutes daily. It takes him 2 minutes per page, and he spends 4 minutes viewing graphics in the book. If x represents the number of pages he reads, then 2x + 4 ≥ 30 represents the time he reads each day. What is the fewest number of pages Fred must read to complete his homework?!!!

OpenStudy (kbreaux9):

help me anyone

OpenStudy (amorfide):

rearrange to solve for x 2x+4>=30 you have 2x+4 the opposite of +4, is -4 so subtract 4 from both sides 2x+4-4>=30-4 2x>=26 now you have 2x, you want x, so the opposite to multiplying by 2, is divide by 2 2x/2>=26/2 x=13

OpenStudy (kbreaux9):

thanks you